Blog Archive

Last modified by Michael Sibayan on 2015/02/13 12:22
Comments (0) · Attachments (0) · History

Blog posts for February 2015

LEO-E Camera track system installation

02/02/2015 - Installation of camera track system begins in East bay. First section in position for assembly after approximately 3 hours.

LEO-C_0_0_0_EX81 registering phantom flow

02/03/2015 - Since the rain event, the Center bay EX81 appears to be continuing to generate data (without flow). Investigation of this anomaly to proceed... 

LEO Flow meter rewiring

EX81 Flow meter re-wiring
02/03/2015 - All 3 EX81 - Tied shield drain wire to EX81 housing and panel earth ground. Data generated around 12:00 are artificial (due to re-wiring).

LEO-E irrigation valve manifold rebuild

02/03/2015 - East bay irrigation valve manifold removed by Tim to perform complete rebuild (07:30).

LEO-E_4_0_1_HMP60 damaged

02/09/2015 - At approximately 08:00 East bay above ground temp/RH probe ( LEO-E_4_0_1_HMP60 ) was damaged by gantry. Probe was torn off it's mount due to insufficient clearance after raising the mast (mast was not raised high enough) and operating the gantry up the hill slope. Data for this probe is (since time of damage) invalid.

81 mm irrigation on LEO-W and LEO-C on 2015-02-12

02/12/2015 - 81 mm Irrigation event
Multi1 circuit on both Center bay and West bay opened for approximately 7 hours 10 minutes.
Start time 09:07
End time 16:18.

Tags:
Created by Michael Sibayan on 2015/02/13 12:22

This wiki is licensed under a Creative Commons 2.0 license
XWiki Enterprise 6.0.1 - Documentation